OpenStudy (smarty):

Simplify (3 + 4i)^2.

OpenStudy (zepp):

Hint: \((a+b)^2=a^2+2ab=b^2\)

OpenStudy (zepp):

\((3+4i)^2=3^2+2(3)(4i)+4i^2\)

OpenStudy (smarty):

I got 9+16i....

OpenStudy (smarty):

A. –7 +24i B. 25 + 24i C. 9 + 16i2 D. –25 + 24i

OpenStudy (anonymous):

FOIL it out... can you simplify from there?